OpenStudy (anonymous):

Determine the slope between the points (-1, -5) and (5, 4) A. 2/3 B. -1/4 C. 1/4 D. 3/2

OpenStudy (anonymous):

Use the slope formula.\[\frac{y_{1} - y_{2}}{x_{1} - x_{2}} = \frac{-5 - 4}{-1 - 5} = ?\]